OpenStudy (anonymous):

Solve for x. 8 – 3x – x = 12

OpenStudy (luigi0210):

Combine like terms: \[\LARGE -(3x+x)+8=12\] What do you get>

OpenStudy (anonymous):

im thinking

OpenStudy (anonymous):

5plus8=12

OpenStudy (luigi0210):

You get \[\LARGE -5x+8=12\] Subtract 8 \[\LARGE -5x=-4\] Divide and finally: \[\LARGE x=\frac{4}{5}\]

OpenStudy (luigi0210):

Sorry, that's suppose to be -4x not -5x.. So x=1, not 4/5